# Fuss-Free Network: A Simplified and Efficient Neural Network for Crowd Counting

### v1

@@ 4.4 Multi-scale fusion experiment
In this section, we conducted comparative experiments on three simple multi-scale fusion strategies designed in Section 3.2. The experiment only adjusted the fusion method of the FFNet and validated it on the Part A and Part B datasets. The experimental results are shown in Table 3 below. In Part A, both the “Add” and “Stepwise addition” methods demonstrate good performance, but the “concatenate” method outperforms them. In Part B, the “concatenate” method shows better performance and robustness. This proves the effectiveness of the “concatenate” method in crowd counting tasks.
@@
Table 3: Results of three simple multi-scale fusion designs on ShanghaiTech Part A and B datasets.
4.5 Ablation Study
As mentioned earlier, our proposed FFNet consists of the following components: a feature extraction backbone, focus transition modules, and a multi-scale fusion structure. We conducted extensive ablation experiments on the ShanghaiTechA dataset to validate the effectiveness of each component in FFNet.
4.5.1 Focus Transition Module
We conducted two ablation experiments on the FTM module, one focusing on local effects and the other on global effects, to verify its effectiveness. First, at the local level, we tested the counting results of the three branches of the model before simple fusion on dataset A. Then, based on this, we removed the FTM module and tested the counting results of the three branches again, and compared the results. The effectiveness of FTM will demonstrate by observing the difference of counting performance before and after the three branches. At the global level, we tested the performance of FFNet with and without FTM to highlight the overall contribution of FTM to FFNet. The results are shown in Table 4. The results indicate that employing the FTM module, whether in local branches or in the overall model, effectively improves the counting accuracy, demonstrating the practicality of the focus transition module.
@@ 4.5.1 Focus Transition Module
Table 4: Effective improvement of model counting performance by FTM.
4.5.2 Loss Function
We also quantitatively evaluated the weighted sum composed of counting loss, variation loss, and optimal transport loss to verify the specific contributions of each loss function to the model’s performance. The results are shown in Table 5. Experimental results show that the model performance reaches its optimal state when the $\lambda_{1}$ and $\lambda_{2}$ are 0.1 and 0.01, respectively. This weighting scheme highlights the central role of count loss in crowd counting prediction, while also moderately considering the image smoothness brought by variation loss and the fine depiction of complex scenes by OT Loss.
Loss function SHA
MAE MSE
$L_{c}$ 55.1 95.3
$L_{c}+L_{ot}+L_{v}$ 55.3 94.9
$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.1)$ 54.6 90.4
$L_{c}+\lambda_{1}L_{ot}+\lambda_{2}L_{v}(\lambda_{2}=0.01)$ 48.3 80.5
Table 5: The total loss function achieved optimal performance when weighting three loss functions and setting $\lambda 1$ and $\lambda 2$ to 0.1 and 0.01, respectively.
4.5.3 Composition Analysis of FFNet
Table 6 displays the component analysis of our model. For the ShanghaiTech PartA dataset, both FTM and the final “Concatenate” fusion contribute significantly to the final MAE metric.
@@ 4.5.3 Composition Analysis of FFNet
Table 6: Composition analysis of FFNet.
4.6 Visualization
In order to further explore the internal mechanism of the model, we visualized the effective receptive field maps and their heat maps of the three branches of the model before simple fusion, so as to reveal the spatial coverage and response intensity distribution of the input data features of each branch of the model before simple fusion. The effective receptive field [39] refers to a pixel on the output feature map of a certain layer in a neural network whose response is influenced by the size of the input image, reflecting the actual sampling range of the layer on the input spatial size. The heatmap is a visualization tool that represents numerical intensity through color coding and is used to display the model’s level of attention or confidence distribution towards specific targets or features in the input data.
@@ 4.6 Visualization
Figure 6: Three branches of the model pass FTM module in the receptive field map and their heat map
Finally, we visualize the model outputs using six images in the NWPU dataset, which exhibit varying scales, densities, lighting conditions and from different scenes. The visualization of the model’s final outputs is presented in Figure 7. Our model performs excellently on images of various scales, demonstrating good technical accuracy.
@@
5 Conclusion
In this paper, we propose a concise and efficient crowd counting model called Fuss-Free Net work (FFNet), which consists only of a backbone network, a multi-scale fusion structure, and focus transition modules. In the concrete implementation, we use the ConvNeXt-Tiny network to build a multi-branch structure, each branch is equipped with a focus transition module, and then the multi-branch features are fused through a simple cascading fusion strategy. Finally, experiments show that the counting accuracy of the model is equal to or even better than that of the complex model, despite its compact structure, small number of parameters, and low computational complexity, breaking the inherent relationship between complexity and performance improvement.
FFNet provides a simple and efficient solution for crowd-counting tasks, but there is still plenty of room to explore. For example, cross-domain adaptability, interpretability, and robustness are improved. In our future work, we will explore the field of crowd counting from the perspectives of simplicity, robustness, and high counting accuracy, hoping to build a more general, reliable, and robust model.

### v3

@@ III-D Multi-scale feature fusion
In this section, we conducted comparative experiments on three simple multi-scale fusion strategies in Figure 4. The experiment only adjusted the fusion method of the FFNet and validated it on the SHA and SHB datasets. The experimental results are shown in Table V below. In SHA, both the “Addition” and “Stepwise addition” operations demonstrate good performance, but the “Concatenate” operation outperforms them. In SHB, the “Concatenate” operation shows better performance and robustness, which proves the effectiveness of the “Concatenate” operation in crowd counting tasks.
@@
TABLE V: Comparison of three simple multi-scale fusion designs on SHA and B datasets.
III-E Ablation Study
As mentioned earlier, our proposed FFNet consists of the following components: a feature extraction backbone, focus transition modules, and a multi-scale fusion structure. We conducted extensive ablation experiments on the SHA dataset to validate the effectiveness of each component in FFNet.
III-E1 Focus Transition Module
@@
TABLE VI: Effective improvement of crowd counting performance through FTM.
We conducted two ablation experiments on the FTM module, one focusing on local effects and the other on global effects, to verify its effectiveness. First, at the local level, we tested the counting results of the three branches of the model before simple fusion on dataset A. Then, based on this, we removed the FTM module and tested the counting results of the three branches again, and compared the results. The effectiveness of FTM will demonstrate by observing the difference of counting performance before and after the three branches. At the global level, we tested the performance of FFNet with and without FTM to highlight the overall contribution of FTM to FFNet. The results are shown in Table VI. The results indicate that employing the FTM module, whether in local branches or in the overall model, effectively improves the counting accuracy, demonstrating the practicality of the focus transition module.
III-E2 Component Analysis of FFNet
@@
TABLE VII: Component analysis of of FFNet.
Table VII displays the component analysis of our model. For the SHA dataset, both FTM and the final “Concatenate” fusion contribute significantly to the final MAE metric.
III-F Effective Receptive Field and Heatmap
Figure 5: Three branches of the model pass FTM module in the receptive field map and their heat map.
To enhance the intuitive understanding of our model’s internal mechanism, we visualized the effective receptive field (ERF) [37] as well as the heat maps of the three branches before fusion. This allows us to uncover the spatial coverage and response intensity distribution of the input data features in each branch before the fusion occurs.
The theoretical Receptive Field (RF) denotes the area in the input space that directly influences the output of an individual neuron in a neural network. However, the contribution of each pixel within a receptive field to the response of an output unit may not be equal. The effective receptive field is employed to quantify the extent to which each pixel on the input feature map affects a pixel on the output feature map. Typically, the impact within a receptive field follows a Gaussian distribution. The heatmap is a visualization tool that represents numerical intensity through color coding and is used to display the model’s level of attention or confidence distribution towards specific targets or features in the input data.
@@ III-F Effective Receptive Field and Heatmap
Finally, we visualize the model outputs using six images in the NWPU dataset, which exhibit varying scales, densities, lighting conditions and from different scenes. The visualization of the model’s final outputs is presented in Figure 6. Our model performs excellently on images of various scales, demonstrating good the fitting ability.
IV Conclusion
In this paper, we propose a simplified and efficient crowd counting model called Fuss-Free Network (FFNet), which consists only of a backbone network, a multi-scale fusion structure with focus transition modules. We use the ConvNeXt-Tiny as backbone network and build a simple feature fusion that contains three branches. Each branch is equipped with its focus transition module, which has the capability to enhance the model’s flexibility and extract a greater variety of features related to the crowd. The features from three branches are fused through a concatenate operation and a $1\times 1$ convolution is applied to generate a single-channel density map. The experiment results show that the counting accuracy of the model is equal to or even better than that of the complex models, despite its compact structure, small number of parameters, and low computational complexity. FFNet provides a simple and efficient solution for crowd counting tasks, but there is still plenty of room to explore. For example, cross-domain adaptability, interpretability, and robustness are improved. In our future work, we will explore the field of crowd counting from the perspectives of simplicity, robustness, and high counting accuracy, hoping to build a more general, reliable, and robust model.